$SPAD/src/input liss2.input

The Axiom Team December 3, 2016

Abstract

Contents

2

— * —

draw(curve(-9*sin(4*t/5),8*sin(t)),t = -5*%pi..5*%pi, [title "Lissajous curve"])

———-

References^{[1] nothing}

4